OpenStudy (anonymous):

If (-1, y) lies on the graph of y = 22x, then y =

OpenStudy (anonymous):

in this orderd pair (-1,y) what is the value of x ?

OpenStudy (anonymous):

@Bonka14 can you answer me ?

OpenStudy (anonymous):

I have no clue thats what I'm trying to figure out..

OpenStudy (anonymous):

okay an ordered pair is always respresented as (x,y) and here you have (-1,y) what is x ?

OpenStudy (anonymous):

Im not sure but my answer choices are -4, 1/4, or 1

OpenStudy (anonymous):

is that all the problem If (-1, y) lies on the graph of y = 22x, then y = or is there something else ?

OpenStudy (anonymous):

Yes that all it says but the problem is If (-1, y) lies on the graph of y = 2^2x, then y = is what? -4 1/4 1

OpenStudy (anonymous):

If just realized on the problem it didn't put the "^" sign.

OpenStudy (anonymous):

okay so I'm gonna help you with this one you have (-1,y) then x=-1 (because an ordered pair is represented as (x,y)) and you have this point (-1,y) lies on the graph of y=2^2*x so all you have to do is to put the value of x in y=2^2*x is it clear ?

OpenStudy (anonymous):

So i multiply the x by the -1?

OpenStudy (anonymous):

no you put instead of x its value

OpenStudy (anonymous):

oh okay so it would be -1

OpenStudy (anonymous):

Right..

OpenStudy (anonymous):

NO wait thats not right.

OpenStudy (anonymous):

so your answer is :

OpenStudy (anonymous):

1? o_O lol im so confused.

OpenStudy (anonymous):

okay slowly (-1,y) ----> x=-1 is it clear til here ?

OpenStudy (anonymous):

Yes so far so good

OpenStudy (anonymous):

Ohh okayy so Would it be -4? Would that be my answer..

OpenStudy (anonymous):

okay, this point (-1,y) belongs to the graph of y=2^2*x so it satisfies it's equation is that clear ?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

good to simplify things, what is a graph ; it contains all the points that satisfies it's equation for example y=2^2*x if x=0 ----> y=2^2*0=0 if x=1-----> y=2^2*1=4*1=4 if x=2-----> y=2^2*2=4*2=8 did you got it ?

OpenStudy (anonymous):

Oh okay yes i did thank you!!(:

OpenStudy (anonymous):

so can you give now the anwser of your question

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

first do you have \[y=2^2x\] or something else?

OpenStudy (anonymous):

@Bonka14

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

what is the expression then can you draw it ?

OpenStudy (anonymous):

is it \[y=(2^2)^x\]